OpenStudy (anonymous):

a square swimming pool ABCD, a man is going to get to C from A, if he is going to walk,the velocity of walking is W,if he is going to swim,the velocity is M,so what he should do to get C with the shortest time?

OpenStudy (anonymous):



OpenStudy (anonymous):

oh i forgot something ( 1.W/M>\[\sqrt{2}\] 2.W/M<\[\sqrt{2}\]

OpenStudy (anonymous):

i need a graph to show how he should do

OpenStudy (anonymous):

\[\min\{\frac{2s}W,\frac{s\sqrt{2}}{M}\}\]

OpenStudy (anonymous):

#1 would make it more effective to walk, while #2 would instead make swimming faster.

OpenStudy (anonymous):

@bahrom7893

OpenStudy (phi):

 let W/M = k or W= k M when is it faster to swim than walk, as a function of k? compare the time to swim to time to walk (replace walking speed with k*M): \[ \frac{s\sqrt{2}}{M} < \frac{2s}{kM} \] solving for k, we find \[ k <\sqrt{2} \] so it is faster to swim when W/M < sqrt(2) otherwise, it is faster to walk